How to find the total electric flux through hemisphere due to a charge placed just above its flat surface

Grade:12

1 Answers

Arun
5 years ago
Dear Mukul

According to gauss law, flux through a closed surface=q/(epsilon).
Here charge enclosed by hemisphere=q/2
therefore, flux =q/2(epsilon)
 
Regards
Arun (askIITians forum expert)